OpenStudy (anonymous):

how would you multiply (x+h+3)(x+3)?

OpenStudy (unklerhaukus):

\[(x+h+3)(x+3)\] \[=x^2+3x+hx+3h+3x+9\]

OpenStudy (anonymous):

You distirbute the terms one by one. So within (x+3), take the x and multiply it by x, then the h, then the 3. Do the same for the 3 from that same binomial (x+3) such that 3 times x, times the h, times the last 3. each of the terms is seaparated by a + sign since they are all positive

OpenStudy (accessdenied):

You can distribute the entire (x+3) to each term of the other expression, like... \[(x+h+3)(x+3) = x(x+3) + h(x+3) + 3(x+3)\] and then distribute into each of those expressions \[x(x+3) + h(x+3) + 3(x+3) = x*x + 3*x + h*x + h*3 + 3*x + 3*3\] then just simplify it down a bit. There are quite a few ways to do it, so this isn't the only one!

OpenStudy (anonymous):

THANKS! :D

OpenStudy (anonymous):

wait, why dont we add up the two 3x's?

OpenStudy (unklerhaukus):

do

OpenStudy (accessdenied):

You can add them together, it's just part of simplification.

OpenStudy (anonymous):

ok